OpenStudy (anonymous):

2x-y=8 7x-3y=27 Using the method of substitution solve these equations. Write an ordered pair.

jimthompson5910 (jim_thompson5910):

Solve the first equation for y 2x-y=8 -y=-2x+8 y = 2x-8 Then plug this into the second equation 7x-3y = 27 7x-3(2x-8) = 27 Keep going to solve for x

jimthompson5910 (jim_thompson5910):

Once you have the solution for x, use this to find the solution for y.

OpenStudy (anonymous):

Do you mind helping me with the steps?

jimthompson5910 (jim_thompson5910):

Sure, the next step is to distribute 7x-3(2x-8) = 27 7x-3(2x)-3(-8) = 27 7x-6x+24 = 27 Then combine like terms x + 24 = 27 What's next?

OpenStudy (anonymous):

take 24 away from both sides?

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

giving you what

OpenStudy (anonymous):

x=3

jimthompson5910 (jim_thompson5910):

now use that to find y

jimthompson5910 (jim_thompson5910):

plug it into y = 2x - 8

OpenStudy (anonymous):

y=2(3)-8 y=6-8 y=2 ( I don't think I did that right)....? that's my best guess

jimthompson5910 (jim_thompson5910):

close

jimthompson5910 (jim_thompson5910):

but 6 - 8 is negative since you're taking a larger number from 6 so 6 - 8 = -2

jimthompson5910 (jim_thompson5910):

making y = -2

OpenStudy (anonymous):

darn, i forgot about the sign....

jimthompson5910 (jim_thompson5910):

think of it in terms of money

OpenStudy (anonymous):

so the ordered pairs are (3, -2) yes?

jimthompson5910 (jim_thompson5910):

you have $6 and you spend $8 you'll be in debt $2 or have -$2 on the balance sheet

jimthompson5910 (jim_thompson5910):

exactly, x = 3 and y = -2 giving you (3,-2) as the solution

OpenStudy (anonymous):

2x-y=8 7x-3y=27 1. take one of these equations 2. bring "x" on one side of equation and rest on other side (so u eliminated x from the other side) 3. put this x = .... any thing in the other equation (substitute that thing on the other side instead of x in the equation) u get an equation in y only 4. solve the equation in y only u get value of y 5. put this value of y in the other equation and solve u got the value of x

OpenStudy (anonymous):

oooh, good idea. Thanks Jim!

jimthompson5910 (jim_thompson5910):

yw